Two people's ages add up to 100. They are 82 years apart. Find the age of each.

I know the answer but we're supposed to set up our work like so:
let x=___
let x+___=___
The thing is I have no idea what the blanks are.
Please help me.
Thank you
P.S.
older=82
yonger=18
I think.

Answer by checkley71(8403)

X+(X+82)=100 OR 2X+82=100 OR 2X=100-82 OR 2X=18 OR X=9 THEN 9+82=91
PROOF 9+91=100 OR 100=100